A rectangle is 4 times as long as it is wide. The area is 196 square feet. What is the perimeter of the rectangle?

Answer by checkley79(3341)

PERIMETER=2L+2W
L=4W
196=2*4W+2W
196=8W+2W
196=10W
196/10=W
19.6=W ANS.
L=4*19.6
L=78.4 ANS.
PROOF:
196=2*78.4+2*19.6
196=156.8+39.2
196=196